OpenStudy (anonymous):

HELP !!!!!!!!!!! Solve for the indicated variable. P = 2L + 2W; for L Solve for the indicated variable. A = 1 + prt; for r

OpenStudy (plainntall):

P = 2L + 2W So you want to express L in terms of P and W correct? We need to start moving everything but L on to the other side of the equation. 2L + 2W= P what happens when we subtract both sides of the equation by 2W?

OpenStudy (anonymous):

there is just 2L=P2W left..... Right????

OpenStudy (plainntall):

Not exactly 2L + 2W= P 2L + 2W-2W= P-2W substract both sides by -2W 2L=P-2W

OpenStudy (plainntall):

You substracted 2W from the left side, but multiplied it on the right

OpenStudy (plainntall):

You always have to do exactly the same thing to both sides of the equation or it messes it all up.

OpenStudy (plainntall):

Now we need to get rid of that 2 on the left side to have just L, what should we do?

OpenStudy (anonymous):

dived 2 on both sides????

OpenStudy (plainntall):

Great! What does that give us then?

OpenStudy (texaschic101):

nice explaining plainntall :)

OpenStudy (anonymous):

L=P-2W/2

OpenStudy (plainntall):

Almost, be careful 2L=P-2W 2L/2=(P-2W)/2 divide both sides by 2 L=(P-2W)/2 or P/2-W

OpenStudy (plainntall):

You have to make sure people know you are dividing P-2W by 2 not just the 2W, that's why I use the parenthesis.

OpenStudy (plainntall):

Now we have to get only r on one side of the equation 1 + prt = A

OpenStudy (anonymous):

should i add 1 to both sides????

OpenStudy (plainntall):

That would give us 1+1+prt = A + 1, we should probably subtract 1 from both sides so the 1 cancels on the left side of the equation

OpenStudy (anonymous):

prt=A+1

OpenStudy (plainntall):

Close, but remember you have to do the same thing on both sides of the equation. You subtracted 1 from the left side, but added 1 to the right. 1-1+prt=A-1 prt=A-1

OpenStudy (plainntall):

What's next to get only r on the left side?

OpenStudy (anonymous):

divide p and t on both sides???

OpenStudy (plainntall):

Tremendous, you got it. Just be careful to use parenthsis.

OpenStudy (anonymous):

ok it would be (A-1)/(pt)

OpenStudy (plainntall):

Terrific, you got that r=(A-1)/(pt) :D

OpenStudy (anonymous):

thankx for the help

OpenStudy (anonymous):

before u go could u help me with one more problem???